# PULL: PU-Learning-based Accurate Link Prediction

### v1

@@ 3.4. Theoretical Analysis
EM (Expectation-Maximization) (Dempster et al., 1977) is an iterative method used for estimating model parameter $\theta$ when there are missing or unobserved data. It assigns latent variables $\mathbf{z}$ to the unobserved data, and maximizes the expectation of the log likelihood $\log p(\mathbf{y},\mathbf{z}\mid\mathbf{X},\theta)$ in terms of $\mathbf{z}$ to optimize $\theta$ where $\mathbf{y}$ and $\mathbf{X}$ are target and input variables, respectively.
@@
In our problem, the target variables are represented as $\mathcal{E}_{\mathcal{P}}$. Thus, the expectation of the log likelihood given the current parameter $\theta$ is written as follows:
(8) $Q(\theta^{\mathrm{new}}\mid\theta)=\mathbb{E}_{\mathbf{z}\sim p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)}[\log p(\mathcal{E}_{\mathcal{P}},\mathbf{z}\mid\mathbf{X},\theta^{\mathrm{new}})]$
@@
where $\theta^{\mathrm{new}}$ is the new parameter. The EM algorithm finds $\theta^{\mathrm{new}}$ that maximizes $Q(\theta^{\mathrm{new}}|\theta)$, and they are used as $\theta$ in the next iteration. The algorithm is widely used in situations involving latent variables as it always improves the likelihood $Q$ through the iterations (Murphy, 2012).
@@
PULL iteratively optimizes the parameter $\theta$ of a link predictor $f$ by minimizing both $\mathcal{L}_{E}^{\prime}$ and $\mathcal{L}_{C}$ where $\mathcal{L}_{E}^{\prime}$ is the approximation of $\mathcal{L}_{E}$ in Equation (5). We compare Equations (5) and (8) to show the similarity between the iterative minimization of $\mathcal{L}_{E}$ in PULL and the iterative maximization of $Q(\theta^{\mathrm{new}}\mid\theta)$ in the EM algorithm.
@@
PULL effectively expresses the distribution $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ in Equation (8) by the multiplication of node potentials in Equation (2). For the joint probability $p(\mathcal{E}_{\mathcal{P}},\mathbf{z}\mid\mathbf{X},\theta^{\mathrm{new}})$ in Equation (8), we approximate it using a link predictor $f_{\theta^{\mathrm{new}}}$ with new parameter $\theta^{\mathrm{new}}$. We consider the link predictor $f_{\theta_{\mathrm{new}}}$ as a marginalization function that gives marginal linking probabilities for each node pair. We also assume that the marginal distributions obtained by $f_{\theta_{\mathrm{new}}}$ are mutually independent. Then the joint probability $p(\mathcal{E}_{\mathcal{P}},\mathbf{z}\mid\mathbf{X},\theta^{\mathrm{new}})$ is approximated as follows:
@@
(9) $\small p(\mathcal{E}_{\mathcal{P}},\mathbf{z}\mid\mathbf{X},\theta^{\mathrm{new}})\approx\prod_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\hat{y}_{ij}\prod_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\bigl{(}z_{ij}\hat{y}_{ij}+(1-z_{ij})(1-\hat{y}_{ij})\bigr{)}$
@@
Using Equations (2) and (9), we derive Theorem 1 that shows the similarity between the iterative minimization of $\mathcal{L}_{E}$ in PULL and the iterative maximization of $Q(\theta^{\mathrm{new}}\mid\theta)$ in the EM algorithm.
@@
Given the assumption in Equation (9), the expected log likelihood $Q(\theta^{\mathrm{new}}\mid\theta)$ of the EM algorithm reduces to the negative of the loss function $\mathcal{L}_{E}$ of PULL with the expected graph $\mathcal{\bar{G}}$:
(10) $\small Q(\theta^{\mathrm{new}}\mid\theta)\approx\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\log\hat{y}_{ij}+\sum_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\bigl{(}\mathbf{A}^{\mathcal{\bar{G}}}_{ij}\log\hat{y}_{ij}+(1-\mathbf{A}^{\mathcal{\bar{G}}}_{ij})\log(1-\hat{y}_{ij})\bigr{)}$
where $\hat{y}_{ij}$ is the estimated linking probability between nodes $i$ and $j$ by $f_{\theta^{\mathrm{new}}}$, and $\mathbf{A}^{\mathcal{\bar{G}}}$ is the corresponding adjacency matrix of $\mathcal{\bar{G}}$ (see Appendix B.1 for proof).
Complexity of PULL
@@
which is linear to the numbers $|\mathcal{V}|$ and $|\mathcal{E}_{\mathcal{P}}|$ of nodes and edges in the given graph $\mathcal{G}_{\mathcal{P}}$, respectively (see Appendix B.2 for proof).
4. Experiments
Table 2. The link prediction accuracy of PULL and baselines in terms of AUROC and AUPRC. Bold numbers denote the best performance, and underlined ones represent the second-best accuracy. PULL outperforms all the baselines in most of the cases.
@@ 4. Experiments
Missing ratio $r_{m}$ = 0.2
Model PubMed Cora-full Chameleon Crocodile Facebook
AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC
[…]
Table 3. Summary of datasets.
Datasets Nodes Edges Features Description
PubMed1 19,717 88,648 500 Citation
Cora-full2 19,793 126,842 8,710 Citation
Chameleon3 2,277 36,101 2,325 Wikipedia
Crocodile3 11,631 191,506 500 Wikipedia
Facebook4 22,470 342,004 128 Social
We conduct extensive experiments on real-world datasets to provide answers to the following questions.
{enumerate*}
Link prediction performance (Section 4.2). How accurate is PULL compared to the baselines for predicting links in edge-incomplete graphs?
Applying PULL to other baselines (Section 4.3). Does applying PULL to other methods improve the link prediction performance?
Effect of iterative learning (Section 4.4). How does the accuracy change over iterations?
Effect of additional loss (Section 4.5). How does the additional loss term $\mathcal{L}_{C}$ of PULL contribute to the performance?
Scalability (Section 4.6). How does the runtime of PULL change as the graph size grows?
4.1. Experimental Settings
Datasets.
We use five real-world datasets from various domains which are summarized in Table 3. PubMed and Cora-full are citation networks where nodes correspond to scientific publications and edges signify citation relationships between the papers. Each node has binary bag-of-words features indicating the presence or absence of specific words from a predefined dictionary. Chameleon and Crocodile are Wikipedia networks, with nodes representing web pages and edges representing hyperlinks between them. Node features include keywords or informative nouns extracted from the Wikipedia pages. Facebook is a social network where nodes represent users, and edges indicate mutual likes. Node features represent user-specific information such as age and gender.
Baselines.
We compare PULL with previous approaches for link prediction in graphs. GCN+CE, GAT+CE, and SAGE+CE use GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2017), and GraphSAGE (Hamilton et al., 2017) for computing linking probabilities, respectively. Cross entropy (CE) loss is utilized for training and $|\mathcal{E}_{\mathcal{P}}|$ non-edges are sampled randomly from $\mathcal{E}_{\mathcal{U}}$ every epoch to balance the ratio between edge and non-edge examples. GAE (Kipf and Welling, 2016b) utilizes an autoencoder to compute the linking probabilities, forcing the predicted graph structure to be similar to the given graph. VGAE (Kipf and Welling, 2016b) exploits a variational autoencoder to learn the embedding of edges based on the given graph structure and node features. ARGA and ARGVA (Pan et al., 2018) respectively improve the performance of GAE and VGAE by introducing adversarial training strategy. VGNAE (Ahn and Kim, 2021) utilizes L2-normalization to obtain better node embeddings for isolated nodes. Bagging-PU (Gan et al., 2022) classifies node pairs into observed and unobserved, and approximates the linking probabilities using the ratio of observed links. More implementation details are described in Appendix C.
Table 4. The performance improvement of baselines with the integration of PULL. The best performance is indicated in bold. Note that PULL enhances the performance of baseline models.

### v2

@@ 3       $\theta\leftarrow\theta^{\mathrm{new}}$;
EM (Expectation-Maximization) (Dempster, Laird, and Rubin 1977) is an iterative method used for estimating model parameter $\theta$ when there are unobserved data. It assigns latent variables $\mathbf{z}$ to the unobserved data, and maximizes the expectation of the log likelihood $\log p(\mathbf{y},\mathbf{z}\mid\mathbf{X},\theta)$ in terms of $\mathbf{z}$ to optimize $\theta$ where $\mathbf{y}$ and $\mathbf{X}$ are target and input variables, respectively.
@@
In our problem, the expectation of the log likelihood given the current parameter $\theta$ is written as follows:
$Q(\theta^{\mathrm{new}}\mid\theta)=\mathbb{E}_{\mathbf{z}\sim p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)}[\log p(\mathcal{E}_{\mathcal{P}},\mathbf{z}\mid\mathbf{X},\theta^{\mathrm{new}})]$ (7)
@@
where $\theta^{\mathrm{new}}$ is the new parameter. The EM algorithm finds $\theta^{\mathrm{new}}$ that maximizes $Q(\theta^{\mathrm{new}}\mid\theta)$, and they are used as $\theta$ in the next iteration. The algorithm is widely used in situations involving latent variables as it always improves the likelihood $Q$ through the iterations (Murphy 2012).
@@
PULL iteratively optimizes the parameter $\theta$ of a link predictor by minimizing both $\mathcal{L}_{E}^{\prime}$ and $\mathcal{L}_{C}$ where $\mathcal{L}_{E}^{\prime}$ is the approximation of $\mathcal{L}_{E}$ in Equation (5). We compare Equations (5) and (7) to show the similarity between the iterative minimization of $\mathcal{L}_{E}$ in PULL and the iterative maximization of $Q(\theta^{\mathrm{new}}\mid\theta)$ in the EM algorithm.
@@
PULL expresses $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ in Equation (7) by the multiplication of node potentials as in Equation (2). For the joint probability $p(\mathcal{E}_{\mathcal{P}},\mathbf{z}\mid\mathbf{X},\theta^{\mathrm{new}})$ in Equation (7), we approximate it using a link predictor $f_{\theta^{\mathrm{new}}}$ with new parameter $\theta^{\mathrm{new}}$. We consider $f_{\theta_{\mathrm{new}}}$ as a marginalization function that gives marginal linking probabilities for each node pair. We also assume that the marginal distributions obtained by $f_{\theta_{\mathrm{new}}}$ are mutually independent. Then the joint probability is approximated as follows:
@@
$p(\mathcal{E}_{\mathcal{P}},\mathbf{z}\mid\mathbf{X},\theta^{\mathrm{new}})\approx\prod_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\hat{y}_{ij}\prod_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\bigl{(}z_{ij}\hat{y}_{ij}+(1-z_{ij})(1-\hat{y}_{ij})\bigr{)}$ (8)
@@
Using Equations (2) and (8), we derive Theorem 1 that shows the similarity between the iterative minimization of $\mathcal{L}_{E}$ in PULL and the iterative maximization of $Q$ in the EM.
@@
Given the assumption in Equation (8), the likelihood $Q(\theta^{\mathrm{new}}\mid\theta)$ of the EM algorithm reduces to the negative of the loss $\mathcal{L}_{E}$ of PULL with the expected graph $\mathcal{\bar{G}}$:
$Q(\theta^{\mathrm{new}}\mid\theta)\approx\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\log\hat{y}_{ij}+\\ \sum_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\bigl{(}\mathbf{A}^{\mathcal{\bar{G}}}_{ij}\log\hat{y}_{ij}+(1-\mathbf{A}^{\mathcal{\bar{G}}}_{ij})\log(1-\hat{y}_{ij})\bigr{)}$ (9)
where $\hat{y}_{ij}$ is the estimated linking probability between nodes $i$ and $j$ by $f_{\theta^{\mathrm{new}}}$, and $\mathbf{A}^{\mathcal{\bar{G}}}$ is the corresponding adjacency matrix of $\mathcal{\bar{G}}$ (see Appendix for proof).
Complexity of PULL.
@@
which is linear to the numbers $|\mathcal{V}|$ and $|\mathcal{E}_{\mathcal{P}}|$ of nodes and edges in $\mathcal{G}_{\mathcal{P}}$, respectively (see Appendix for proof).
Experiments
Table 1: The link prediction accuracy of PULL and baselines in terms of AUROC and AUPRC. Bold numbers denote the best performance, and underlined ones represent the second-best accuracy. PULL outperforms all the baselines in most of the cases.
@@
Table 2: The performance improvement of baselines with the integration of PULL. The best performance is indicated in bold. Note that PULL enhances the performance of baseline models.